OpenStudy (anonymous):

solving systems of linear equations in three variables- ok i dont know how to do this! I need help! heres the problem: 2x+y+z=-10 -x-4y+z=12 2x+2y+3z=13 thank you so much!

OpenStudy (campbell_st):

well you have 3 equations in 3 unknowns, and you want to reduce to to 2 equations in 2 unknowns label the equations top = 1 middle = 2 and bottom =3 the easiest elimination is x so equation 2 times 2 + equation 1 2x + y + z = -10 -2x - 8y + 2z = 24 ----------------- -7y + 3z = 14 this is equation 4 then equation 2 times 2 + equation 3 -2x -8x + 2z = 24 2x + 2y + 3z = 13 ---------------- -6y + 5z = 37 equation 5 now you have 2 equations in 2 unknowns.... so eliminate z. equation 4 times 5 + equation 5 times - 3 -35y + 15z = 70 18y -15z = -111 --------------- -17y = -41 so you can now solve for y. then when you get it subsitute into either equation 4 or 5 to find z when go to equation 1 and substiute y and z to find x. I can tell you the solution contains ugly fractions. hope it helps